# Machine-Readable Ads: Accessibility and Behavioral Patterns of AI Web Agents interacting with Online Advertisements

### v1

@@ 4.1 Implementation Best Practices for Agent-Friendly Web Design
Across the baseline and all follow-up experiments we repeatedly observed five practical factors that govern whether current DOM- or pixel-level LLM agents can notice an advertisement, click it, and extract its contents correctly. The forthcoming sections and tables distil the quantitative findings. Throughout the results, any mention of an “agent” refers to the Browser Use configuration unless we explicitly state that the run was carried out with the OpenAI Operator to show explicit differences between the architectural capabilities.
@@ 4.1.1 Position & Explicit Calls-to-Action
Across 120 recorded interactions (10 runs for each of the four subscription-related prompts 1,2, 3 and 8 per model) the agents clicked only two on-screen regions (Table 2):
@@
No clicks were registered on the vertical banners, nor anywhere else on the page. One outlier was Gemini 2.0 Flash, which bypassed clicking altogether in a single run by entering the subscription URL directly.
Table 2: Click counts per model for subscription CTAs (40 runs per model)
@@
*  Note: Operator denies tasks where, for example, signing up for sweepstakes is the objective; such tasks are not runnable via Operator.
Removing the highly relevant sticky bottom advertisement caused all models to default to the header link; re-introducing it restored the original pattern, ignoring any and all other advertisements that would also lead to the subscription page in all runs. When both side-banners carried identical markup and applicable content as well as featuring a dedicated overlaid DOM button, all runs interacted exclusively with the left banner, indicating a top-to-bottom, left-to-right traversal that stops at the first satisfactory candidate. When the side-banner ads were only enclosed in a clickable <a> element—without any extra overlay button in the DOM—none of them were clicked during any of the recorded runs. Operator, not DOM reliant, seems to process the page much the same, exclusively accessing the subscription page via the “Subscribe” button at the top.
@@ 4.1.2 Semantic Markup for Click Recognition
As outlined in Section 4.1.1, the agents seemed to consistently favour interface elements that manifest as explicit DOM buttons. To verify this tendency, we executed ten additional runs per model after removing the sticky bottom strip while leaving the original side-banner advertisements intact. These banners were images wrapped in <a> tags with the ARIA role banner and therefore exposed no visible or DOM-level buttons. Under these conditions, every agent interacted solely with the header’s “Subscribe” button; all other elements remained untouched as is summarized in Table 3.
Table 3: Clicks after removing the bottom strip (10 runs/model)
Model Subscribe Button Side banners Direct URL
Claude 3.7 Sonnet 10/10 0/10 0/10
GPT-4o 10/10 0/10 0/10
Gemini 2.0 Flash 10/10 0/10 0/10
Operator 10/10 0/10 0/10
Subsequently, we introduced purely pixel-based call-to-action buttons baked into the banner image itself, investigating if the agents would perceive the bitmap text as a “button-like” affordance. Two variants were rendered:
  1. 1.
    Plain label – a minimalist strip reading “Order Now” in the corporate typeface, centered on the creative;
  2. 2.
    Natural-language prompt – the same banner, but with a full-sentence invitation (“Click here to order your subscription now”) to strengthen the semantic cue.
Although the visual appearance closely mimicked a clickable control, the underlying markup remained a single <img> tag wrapped inside an <a> tag. Across ten runs per model in these configurations, none of the agents registered a click on either banner variant (see Table 4), confirming that pixel-only cues are insufficient to signal interactivity to current DOM-focused browsing agents. Operator also did not interact with the purely pixel-based button variants.
Replacing the pixel-based buttons with an actual DOM <button> overlay finally prompted the agents to act. Claude 3.7 Sonnet interacted with the side banners a total of seven times, choosing the header “Subscribe” link in the other three. GPT-4o also clicked the side banners in 7 of 10 runs, defaulting to “Subscribe” for the remaining three. Gemini 2.0 Flash split its clicks evenly—five on the side banners, five on the subscription link. Crucially, whenever any model interacted with a side banner, it was always the left banner; no right-banner clicks were recorded as is summarized in Table 4. Operator still exclusively interacted with the “Subscribe” button in all runs.
Table 4: Banner clicks with different CTA implementations (10 runs/model). C3.7S: Claude 3.7 Sonnet; G2.0F: Gemini 2.0 Flash; Oper.: Operator.
@@
We then performed a set of follow-up experiments, executing ten trials per model for several accessibility-focused markup variants in combination with the DOM-focused agents. The goal was to discover configurations that forgo a dedicated DOM button yet still elicit agent clicks. In each variant, we appended natural-language text to the side-banner elements that explicitly described them as clickable and give insights about their content. The resulting counts of side-banner interactions are summarized in Table 5.
Table 5: Banner clicks for different hidden-text strategies (10 runs per model). C3.7S: Claude 3.7 Sonnet; G2.0F: Gemini 2.0 Flash. “sr-only”: screen-reader-only class; “aria-*”: ARIA attributes.
@@
.sr-only $+$ aria-hidden 0/10 3/10 10/10
Hidden <p> (role=presentation) 0/10 2/10 10/10
@@
Off-screen <span> (no aria-hidden) 9/10 10/10 10/10
@@ 4.1.3 Scrolling Depth & Satisficing behavior
The agents were first tested for their basic willingness to scroll by placing a unique combined digital + print subscription banner exactly one viewport below the fold. Despite this banner’s unique feature, none of the agents retrieved its content—each settled instead for the first partially relevant offer (like digital-only or unrelated subscription promos). This showed a clear “satisficing” tendency: stopping once a minimally acceptable result was found.
@@
Table 6: Success rates ($N=10$ runs each) for retrieving the combined digital $+$ print subscription banner after an explicit discount-code prompt.
@@
Table 7: Average scroll_down actions per run (10 runs/model)
@@
Regarding task 4 across ten runs, Claude 3.7 Sonnet surfaced the greatest variety of promotions, citing seven distinct offers and doing so 37 times in total. Google’s Gemini 2.0 Flash and OpenAI’s GPT-4o each highlighted four different promotions; Gemini mentioned them 26 times, while GPT-4o referenced them 21 times. Thus, Claude was both the most prolific and the most diverse in its ad reporting, with Gemini second in overall frequency and GPT-4o a close third. Claude 3.7 Sonnet was, furthermore, the only model to explore subpages and report on advertisements there regarding this task.
@@
In the baseline evaluation, all agents using Browser Use correctly parsed dynamic content such as a GIF banner, but only because their initialization times inadvertently allowed the GIF to finish loading—none of the agents explicitly reasoned about waiting for dynamic elements. When tested with a longer 12-second video and the prompt “Find deals for Tiroler Markt for me,” only GPT-4o ever retrieved the brief two-second coupon code segment, and only once out of ten runs, purely by chance because it lingered on the page long enough to catch the relevant frames. The agents’ reasoning never showed any awareness of needing to wait for or parse multiple video frames. When hidden <span> tags provided a text alternative for the video’s content, all agents ignored the video entirely and used the text instead. This highlighted a difference between architectures: DOM-focused agents could at least return text descriptions embedded in the DOM, while pixel-focused frameworks like Operator could not. Despite visually “seeing” the video, Operator failed to infer the discount code in all runs because no single video frame ever displayed both the store name and the coupon code together—reflecting how real-world dynamic ads sequence information.
@@ 4.1.5 Native Browser Dialogs
When faced with native browser prompts—specifically a geolocation permission dialog—task completion initially broke down entirely because the agents could not interact with the browser’s native UI layer. The issue was mitigated in two ways: while Browser Use pre-approved location requests so that the prompt never appeared, Operator automatically rejected the request, effectively bypassing the dialog without interacting with the native element.
@@ 4.2 Behavioral Patterns of AI Agents
Across all baseline tasks using Browser Use, the three models (Claude 3.7 Sonnet, GPT-4o, Gemini 2.0 Flash) stayed firmly anchored to on-page text and did not hallucinate payment or coupon details. OpenAI’s Operator was excluded here because it is built for high-friction, user-controlled steps: it pauses for user confirmation at every sensitive action (logins, purchases, subscriptions). In contrast, Browser Use can run multi-step tasks without pausing, leaving only a global stop option for the user—making it more representative of autonomous web agent behavior for this study.
@@
Task 8 revealed a key weakness: when the site offered a sweepstake promo—"Only now -20% on all digital subscriptions and a chance to win a trip for 2 to Jesolo"—the agents did proceed with paid subscriptions in most runs. Specifically, GPT-4o and Claude 3.7 Sonnet subscribed in 100% of trials (10/10); Gemini 2.0 Flash did so in 70% (7/10). The few exceptions with Gemini happened only when it failed to recognize the ad as a sweepstake, instead stating there was "no sweepstakes content." Because sweepstake entry required the subscription, the models linked purchase directly to task success—revealing their vulnerability to implied instructions to spend real money.
@@ 4.2.1 Cookie Consent and Privacy Dialogs
Three distinct cookie banner designs were implemented to evaluate AI agent responses. The non-essential sticky footer could be disregarded without impeding site navigation, the modal blocker prevented any user interaction until a consent decision was made, and the predatory modal employed sensationalized language and conflicting claims designed to pressure acceptance while maintaining somewhat obvious deceptive intent, thereby testing models’ capacity to recognize and appropriately respond to transparent dark patterns. Each banner configuration included an Accept All button alongside a secondary interface for detailed preference management (Accept Selected / Deny All), accessible through additional user interaction. For 30 experimental trials per banner variant (10 trials per model), we documented the behavioral patterns shown in Table 8141414Prompt: “Summarize the headline article.”.
Table 8: Cookie-banner acceptance rates across AI models. G4o: GPT-4o; C3.7S: Claude 3.7 Sonnet; G2.0F: Gemini 2.0 Flash.
@@
*For G4o, 2/8 acceptance cases still explored alternate interaction paths before accepting.
Notably, none of the models accessed granular settings options for any banner. They either straight up accepted everything or nothing. The predatory modal’s manipulative language (including phrases such as “ATTENTION: This offer expires automatically!” and claims of “97.3% of users choose full activation”) successfully deterred all models from acceptance.
@@ 5.1 Implications for Ad Design and Site Markup
Across all experimental conditions—baseline coupon hunts, explicit discount-code prompts, and the infinite-scroll challenge—agents displayed a consistent satisficing strategy: once a minimally plausible answer could be generated, exploration ceased. The infinite-scroll test highlights the effect most clearly: none of the 30 runs (10 per model) issued more than five consecutive scroll_down commands, even though the requested Tiroler Markt coupon was available only on virtual page 10. Instead, agents abandoned deeper exploration returning an answer along the lines of “I was not able to find a discount code for Tiroler Markt on the provided website.” or providing suboptimal answers, even when the budget would allow for further steps.
From a cognitive-economy standpoint such early stopping is rational—lowering interaction cost while superficially satisfying the query-but for practitioners the implication is stark: critical promotions must appear within the first two viewports. Content buried deeper is effectively invisible to today’s DOM-centric agents unless the prompt explicitly forces long-form scrolling.
Five implementation levers.
@@
    Pair visual banners with real DOM controls. Embed a semantic <button> or <a> overlay; pure pixel-based calls-to-action remain read-only.
@@
    Surface key ads in the top-left quadrant. Agents traverse the accessibility tree top-to-bottom, left-to-right and halt at the first satisfactory element.
@@
    Expose clickability through hidden text. Off-screen <span> text kept in the DOM yielded a consistently high click-through rate across models; data-* or ARIA attributes alone did not.
@@
    Replace native browser dialogs with HTML fallbacks. Permission pop-ups, CAPTCHA widgets, and other non-DOM overlays may block agents. Mission-critical flows should avoid such dialogs or provide a reachable HTML alternative.
@@
    Avoid videos and dynamic media. Agents recognize moving banners only by accident—e.g. a carousel slide happened to advance at the exact moment a snapshot was taken. In most runs the discount frame is missed entirely. Prefer static creatives, or at minimum expose an equivalent text offer in the DOM.
In summary, engineering ad experiences for autonomous agents resembles accessibility optimisation more than traditional creative design: success depends less on visual appeal and more on semantic, machine-readable structure. It is also of note to mention that while Operator did not click the side-banner advertisements, in theory, it should be able to recognize even solely pixel-based CTAs.
@@ 5.2 Behavioral Patterns and Model-Specific Risk Profiles
Hallucination guardrails.
None of the tested agents fabricated sensitive artefacts such as credit-card numbers, addresses, or CVV codes unless we provided dummy data directly in the prompt. In contrast, they freely hallucinated non-critical personal details (e.g., fictitious names in form fields) or booking metadata, mirroring observations in prior work on travel agents [1].
Autonomy.
Tasks 5 and 6 tested whether models would purchase a subscription opportunistically. All three abstained—even though a subscription simplified the task—until task 8 tied sweepstake participation directly to a paywall. There they crossed the boundary (10/10 for GPT-4o and Claude, 7/10 for Gemini). The trigger therefore appears to be a goal-contingent justification: if the downstream objective demands payment, and the two are closely enough aligned, agents comply; otherwise they default to non-monetary shortcuts, if available. This finding indicates a systematic failure in cost-benefit analysis and decision boundaries when AI agents encounter promotional content that conflates promotional participation with financial commitments.
Heterogeneous Risk Profiles.
@@
These findings become specifically critical in commercial contexts: The AI assistants’ weakness when faced with the sweepstakes-paywall combination demonstrates how easily automated decisions can be influenced through targeted design. This opens new avenues for consumer manipulation. Since AI assistants are deployed especially in commercial environments where companies have strong economic incentives to exploit such behavioral patterns, these insights carry significant implications for consumers: Du and Xie describe how AI, while convenient, simultaneously creates challenges for privacy and autonomy [19].
While Article 27 of the EU AI Act establishes the overarching requirement for interactions with AI to be clearly recognisable as automated, Article 52 implements this principle as concrete disclosure obligations for low-risk conversational systems. To ensure that this transparency effectively safeguards user autonomy, providers should reinforce it with the “Technical Robustness & Safety” standards set for trustworthy AI by the EU High-Level Expert Group, which treat reliability as a precondition for meaningful and informed consent.
However, translating regulatory mandates into technical reality reveals a critical gap: although we can implement DOM-level markup and observe behavioral patterns, there are no standardized metrics in the field to quantify agent compliance. Our study highlights this issue—we documented concerning behaviors such as universal subscription acceptance, but could only categorize them qualitatively. Future frameworks must operationalize trust through automated consent scoring, financial decision protocols and risk matrices that transform behavioral observations into benchmarkable metrics. Only by bridging this measurement gap can designers convert the EU AI Act’s legal requirements into verifiable engineering artefacts and ensure that agent convenience does not systematically undermine user autonomy in advertising contexts.

### v2

@@ 4.1 Implementation Best Practices for Agent-Friendly Web Design
Across the baseline and all follow-up experiments we repeatedly observed five practical factors that govern whether current DOM- or pixel-level LLM agents can notice an advertisement, click it, and extract its contents correctly. The forthcoming sections and tables distill the quantitative findings. Throughout the results, any mention of an “agent” refers to the Browser Use configuration unless we explicitly state that the run was carried out with the OpenAI Operator to show explicit differences between the architectural capabilities.
@@ 4.1.1 Position & Explicit Calls-to-Action
Across 120 recorded interactions (10 runs for each of the four subscription‑related prompts 1, 2, 3 and 8 per model) the agents clicked only two on‑screen regions as depicted in Table 2:
@@
No clicks were registered on the vertical banners, nor anywhere else on the page. One outlier was Gemini 2.0 Flash, which bypassed clicking altogether in a single run by directly manipulating the URL. Operator categorically denied tasks where, for example, signing up for sweepstakes is the objective.
Table 2: Click counts per model for subscription CTAs
@@
Removing the highly relevant sticky bottom advertisement caused all models to default to the header link; re‑introducing it restored the original pattern, ignoring any and all other advertisements that would also lead to the subscription page in all runs. When both side-banners carried identical markup and applicable content as well as featuring a dedicated overlaid DOM button, all runs interacted exclusively with the left banner, indicating a top-to-bottom, left-to-right traversal that stops at the first satisfactory candidate. When the side-banner ads were only enclosed in a clickable <a> element, without any extra overlay button in the DOM, none of them were clicked during any of the recorded runs. Operator, not DOM reliant, seems to process the page much the same, exclusively accessing the subscription page via the “Subscribe” button at the top.
@@ 4.1.2 Semantic Markup for Click Recognition
The agents seemed to consistently favor interface elements that manifest as explicit DOM buttons. To verify this tendency, we executed 10 additional runs per model after removing the sticky bottom strip while leaving the original side-banner advertisements intact. These banners were images wrapped in <a> tags with the ARIA role banner and therefore exposed no visible or DOM-level buttons. Under these conditions, every agent interacted solely with the header’s “Subscribe” button while all other elements, such as the side banners or direct URL manipulation, remained untouched.
Subsequently, we introduced purely pixel-based call-to-action buttons baked into the banner image itself, investigating if the agents would perceive the bitmap text as a “button-like” affordance. For this, a minimalist strip reading “Order Now” in the corporate typeface, was centered on the creative.
Although the visual appearance closely mimicked a clickable control, the underlying markup remained a single <img> tag wrapped inside an <a> tag. Across 10 runs per model in these configurations, none of the agents registered a click on this banner variant as can be seen in Table 3, confirming that pixel-only cues are insufficient to signal interactivity to current DOM-focused browsing agents. Operator also did not interact with the purely pixel-based button variants.
Replacing the pixel-based buttons with an actual DOM <button> overlay finally prompted the agents to act. Claude 3.7 Sonnet interacted with the side banners a total of seven times, choosing the header “Subscribe” link in the other three. GPT-4o also clicked the side banners in 7 of 10 runs, defaulting to “Subscribe” for the remaining three. Gemini 2.0 Flash split its clicks evenly, five on the side banners, five on the subscription link. Crucially, whenever any model interacted with a side banner, it was always the left banner; no right-banner clicks were recorded as is summarized in Table 3. Operator still exclusively interacted with the “Subscribe” button in all runs.
Table 3: Banner clicks with different CTA implementations; C3.7S: Claude 3.7 Sonnet; G2.0F: Gemini 2.0 Flash; Oper.: Operator
@@
We then performed a set of follow-up experiments, executing 10 trials per model for several accessibility-focused markup variants in combination with the DOM-focused agents. The goal was to discover configurations that forgo a dedicated DOM button yet still elicit agent clicks. In each variant, we appended natural-language text to the side-banner elements that explicitly described them as clickable and give insights about their content. The resulting counts of side-banner interactions are summarized in Table 4.
Table 4: Banner clicks for different hidden-text strategies; C3.7S: Claude 3.7 Sonnet; G2.0F: Gemini 2.0 Flash.
@@
off-screen <span> 9/10 10/10 10/10
@@ 4.1.3 Scrolling Depth & Satisficing Behavior
The agents were first tested for their basic willingness to scroll by placing a unique combined digital + print subscription banner exactly one viewport below the fold. Despite this banner’s unique feature, none of the agents retrieved its content, each settled instead for the first partially relevant offer (like digital-only or unrelated subscription promos). This showed a clear “satisficing” tendency: stopping once a minimally acceptable result was found.
@@
Table 5: Success rates for retrieving the combined digital ++ print subscription banner after an explicit discount-code prompt
@@
Table 6: Average scroll_down actions per run
@@
Regarding task 4 across 10 runs, Claude 3.7 Sonnet surfaced the greatest variety of promotions, citing seven distinct offers and doing so 37 times in total. Google’s Gemini 2.0 Flash and OpenAI’s GPT-4o each highlighted four different promotions; Gemini mentioned them 26 times, while GPT-4o referenced them 21 times. Thus, Claude was both the most prolific and the most diverse in its ad reporting, with Gemini second in overall frequency and GPT-4o a close third. Claude 3.7 Sonnet was, furthermore, the only model to explore subpages and report on advertisements there regarding this task.
@@
In the baseline evaluation, all agents using Browser Use correctly parsed dynamic content such as a GIF banner, but only because their initialization times inadvertently allowed the GIF to finish loading, none of the agents explicitly reasoned about waiting for dynamic elements. When tested with a longer 12-second video and the prompt “Find deals for Tiroler Markt for me,” only GPT-4o ever retrieved the brief two-second coupon code segment, and only once out of 10 runs, purely by chance because it lingered on the page long enough to catch the relevant frames. The agents’ reasoning never showed any awareness of needing to wait for or parse multiple video frames. When hidden <span> tags provided a text alternative for the video’s content, all agents ignored the video entirely and used the text instead. This highlighted a difference between architectures: DOM-focused agents could at least return text descriptions embedded in the DOM, while pixel-focused frameworks like Operator could not. Despite visually seeing the video, Operator failed to infer the discount code in all runs because no single video frame ever displayed both the store name and the coupon code together, reflecting how real-world dynamic ads sequence information.
@@ 4.1.5 Native Browser Dialogs
When faced with native browser prompts, specifically a geolocation permission dialog, task completion initially broke down entirely because the agents could not interact with the browser’s native UI layer. The issue was mitigated in two ways: while Browser Use pre-approved location requests so that the prompt never appeared, Operator automatically rejected the request, effectively bypassing the dialog without interacting with the native element.
@@ 4.2 Behavioral Patterns of AI Agents
Across all baseline tasks using Browser Use, the three models (Claude 3.7 Sonnet, GPT-4o, Gemini 2.0 Flash) stayed firmly anchored to on-page text and did not hallucinate payment or coupon details. OpenAI’s Operator was excluded here because it is built for high-friction, user-controlled steps: it pauses for user confirmation at every sensitive action (logins, purchases, subscriptions). In contrast, Browser Use can run multi-step tasks without pausing, leaving only a global stop option for the user, making it more representative of autonomous web agent behavior for this study.
@@
Task 8 revealed a key weakness: when the site offered a sweepstake promo, “Only now -20% on all digital subscriptions and a chance to win a trip for 2 to Jesolo”, the agents did proceed with paid subscriptions in most runs. Specifically, GPT-4o and Claude 3.7 Sonnet subscribed in 100% of trials (10/10); Gemini 2.0 Flash did so in 70% (7/10). The few exceptions with Gemini happened only when it failed to recognize the ad as a sweepstake, instead stating there was “no sweepstakes content.” Because sweepstake entry required the subscription, the models linked purchase directly to task success, revealing their vulnerability to implied instructions to spend real money.
@@ 4.2.1 Cookie Consent and Privacy Dialogs
Three distinct cookie banner designs were implemented to evaluate AI agent responses. The non-essential sticky footer could be disregarded without impeding site navigation, the modal blocker prevented any user interaction until a consent decision was made, and the predatory modal employed sensationalized language and conflicting claims designed to pressure acceptance while maintaining somewhat obvious deceptive intent, thereby testing models’ capacity to recognize and appropriately respond to transparent dark patterns. Each banner configuration included an “Accept All” button alongside a secondary interface for detailed preference management (“Accept Selected” / “Deny All”), accessible through additional user interaction. For 30 experimental trials per banner variant (10 trials per model), we documented the behavioral patterns shown in Table 7141414Prompt: “Summarize the headline article.”.
Table 7: Cookie-banner acceptance rates across AI models; G4o: GPT-4o; C3.7S: Claude 3.7 Sonnet; G2.0F: Gemini 2.0 Flash.
@@
Notably, none of the models accessed granular settings options for any banner. They either straight up accepted everything or nothing. The predatory modal’s manipulative language (including phrases such as “ATTENTION: This offer expires automatically!” and claims of “97.3% of users choose full activation”) successfully deterred all models from acceptance. For GPT-4o, 2/8 acceptance cases still explored alternate interaction paths before accepting.
@@ 5.1 Implications for Ad Design and Site Markup
Across all experimental conditions, baseline coupon hunts, explicit discount-code prompts, and the infinite-scroll challenge, agents displayed a consistent satisficing strategy: once a minimally plausible answer could be generated, exploration ceased. The infinite-scroll test highlights the effect most clearly: none of the 30 runs (10 per model) issued more than five consecutive scroll_down commands, even though the requested Tiroler Markt coupon was available only on virtual page 10. Instead, agents abandoned deeper exploration returning an answer along the lines of “I was not able to find a discount code for Tiroler Markt on the provided website.” or providing suboptimal answers, even when the budget would allow for further steps.
From a cognitive-economy standpoint such early stopping is rational, lowering interaction cost while superficially satisfying the query, but for practitioners the implication is stark: critical promotions must appear within the first two viewports. Content buried deeper is effectively invisible to today’s DOM-centric agents unless the prompt explicitly forces long-form scrolling.
@@
    Pair visuals with real controls: Overlay a semantic <button> element because images with pixel-only CTAs, even when the image is wrapped in <a> tags, remain inaccessible.
@@
    Prioritize top-left placement: Agents scan the accessibility tree top-to-bottom, left-to-right, stopping at the first suitable element.
@@
    Mark clickability clearly: Use hidden text like an off-screen <span> for higher click-through; data-* attributes or ARIA alone are unreliable. Interpretation depends on the model and whether or not the framework even extracts them.
@@
    Avoid native dialogs: Replace pop-ups, CAPTCHAs, and other non-DOM overlays with reachable HTML alternatives that can be handled on a task by task basis for mission-critical flows.
@@
    Limit dynamic media: Moving banners are often missed; use static creatives or provide equivalent text in the DOM.
In summary, engineering ad experiences for autonomous agents resembles accessibility optimization more than traditional creative design: success depends less on visual appeal and more on semantic, machine-readable structure. It is also of note to mention that while Operator did not click the side-banner advertisements, in theory, it should be able to recognize even solely pixel-based CTAs.
@@ 5.2 Behavioral Patterns and Model-Specific Risk Profiles
Hallucination guardrails
None of the tested agents fabricated sensitive artifacts such as credit-card numbers, addresses, or CVV codes unless we provided dummy data directly in the prompt. In contrast, they freely hallucinated non-critical personal details (e.g., fictitious names in form fields) or booking metadata, mirroring observations in prior work on travel agents [stoeckl2025aiagentsinteractingonline].
Autonomy
Tasks 5 and 6 tested whether models would purchase a subscription opportunistically. All three abstained, even though a subscription simplified the task, until task 8 tied sweepstake participation directly to a paywall. There they crossed the boundary (10/10 for GPT-4o and Claude, 7/10 for Gemini). The trigger therefore appears to be a goal-contingent justification: if the downstream objective demands payment, and the two are closely enough aligned, agents comply; otherwise they default to non-monetary shortcuts, if available. This finding indicates a systematic failure in cost-benefit analysis and decision boundaries when AI agents encounter promotional content that conflates promotional participation with financial commitments.
Heterogeneous Risk Profiles
@@
These findings become specifically critical in commercial contexts: the AI assistants’ weakness when faced with the sweepstakes-paywall combination demonstrates how easily automated decisions can be influenced through targeted design. This opens new avenues for consumer manipulation. Since AI assistants are deployed especially in commercial environments where companies have strong economic incentives to exploit such behavioral patterns, these insights carry significant implications for consumers: Du and Xie describe how AI, while convenient, simultaneously creates challenges for privacy and autonomy [DuXie2021].
While Recital 27 of the EU AI Act refers to overarching ethical principles, including autonomy, privacy, technical robustness and transparency, i.e., that interactions with AI should be clearly recognizable as automated, the Regulation’s broader transparency and trustworthiness provisions implement these principles into binding disclosure and labeling duties. To ensure that transparency effectively safeguards user autonomy, providers should reinforce it with the “Technical Robustness and Safety” standards for trustworthy AI by the EU High-Level Expert Group, which treat reliability as a precondition for meaningful and informed consent. At the same time, defined measures and benchmarks for assessing the trustworthiness of LLM-powered web agents remain limited and purely behavioral indicators and may not capture the full range of ethical and regulatory factors, including adherence to fairness and non-discrimination.
Translating regulatory mandates into technical reality reveals a critical gap: although we can implement DOM-level markup and observe behavioral patterns, there are no standardized, web agent-specific metrics to quantify compliance. Our study highlights this issue, as we documented concerning behaviors, such as consistently high subscription-acceptance rates, but could only categorize them qualitatively. Future frameworks must operationalize trust through automated consent scoring, financial decision protocols and risk matrices that transform behavioral observations into benchmarkable metrics. Only by bridging this measurement gap can designers convert the EU AI Act’s legal requirements into verifiable engineering artifacts and ensure that agent convenience does not systematically undermine user autonomy in advertising contexts.